# MERLIN: Multimodal Embedding Refinement via LLM-based Iterative Navigation for Text-Video Retrieval-Rerank Pipeline

### v1

@@ Abstract
The rapid expansion of multimedia content has made accurately retrieving relevant videos from large collections increasingly challenging. Recent advancements in text-video retrieval have focused on cross-modal interactions, large-scale foundation model training, and probabilistic modeling, yet often neglect the crucial user perspective, leading to discrepancies between user queries and the content retrieved. To address this, we introduce MERLIN (Multimodal Embedding Refinement via LLM-based Iterative Navigation), a novel, training-free pipeline that leverages Large Language Models (LLMs) for iterative feedback learning. MERLIN refines query embeddings from a user perspective, enhancing alignment between queries and video content through a dynamic question answering process. Experimental results on datasets like MSR-VTT, MSVD, and ActivityNet demonstrate that MERLIN substantially improves Recall@1, outperforming existing systems and confirming the benefits of integrating LLMs into multimodal retrieval systems for more responsive and context-aware multimedia retrieval [^1] Code will be released upon acceptance..
@@
Donghoon Han1111, Eunhwan Park1111, Gisang Lee2111, Adam Lee3, 4, Nojun Kwak5222 1Buzzni AI Lab, 2KAIST, 3UC Berkeley, 4Wrtn, 5Seoul National University {owen, jude}@buzzni.com, bobopack@kaist.ac.kr, alee00@berkeley.edu, nojunk@snu.ac.kr
footnotetext: 111 Equal contribution.footnotetext: 222 Corresponding author.
@@ 1 Introduction
Multimedia content has recently grown rapidly in both quantity and quality, making the task of finding relevant videos from vast collections increasingly challenging. While recent studies on text-video retrieval have primarily focused on cross-modal interaction Wang et al. (2023); Huang et al. (2023); Wu et al. (2023); Jin et al. (2023), large-scale foundation model training Chen et al. (2024b, 2023); Zhao et al. (2024); Wang et al. (2024a) and probabilistic modeling Hao et al. (2023); Fang et al. (2023); Hao and Zhang (2023), there remains a notable lack of consideration for the discrepancy in text-video retrieval. For instance, as illustrated in Figure 1, the video caption “a baby playing with a cat’s tail” fails to fully capture the additional context of a playful interaction between the baby and the cat. In real-world scenarios, such discrepancies often arise because users tend to submit succinct queries that do not capture the full context of the videos related to their search intent. Consequently, this mismatch can lead to unsatisfactory retrieval performance. Moreover, neglecting the user perspective, making users to refine their natural language query for multiple times to fully reflect their search intent, degrades the quality of user experience and makes it difficult to understand the search intent, leading to a discrepancy between user queries and the information within the retrieved videos.
@@
Figure 1: An illustration of the discrepancy between the video caption which could be treated as a user query and the video from MSR-VTT dataset. blue and red indicate the additional details within the video frame, and those reflecting temporal information, respectively.
@@
To address this issue, we introduce MERLIN (Multimodal Embedding Refinement via LLM-based Iterative Navigation), a novel training-free and iterative feedback learning pipeline that leverages the power of Large Language Models (LLMs) to augment queries based on the user perspective, thereby mitigating the aforementioned discrepancies and significantly improving the text-video retrieval performance. Inspired by human problem-solving and cognitive feedback mechanisms Flower and Hayes (1981); Doherty and Balzer (1988), we employ an interactive and iterative feedback learning Böhm et al. (2019); Stiennon et al. (2020); Ziegler et al. (2019); Wu et al. (2020); Ouyang et al. (2022); Glaese et al. (2022); Akyürek et al. (2023); Madaan et al. (2023); Lee et al. (2024); Liang et al. (2024) consisting of a question-and-answer process that iteratively refines query embeddings for text-video retrieval. Moreover, to our best knowledge, MERLIN presents the first implementation of a retrieval–rerank pipeline in the domain of text-video retrieval, establishing a novel framework that prioritizes user intention and interaction in refining search results.
The primary strength of MERLIN lies in its capability to iteratively adapt and refine query embeddings without necessitating the costly re-training of pre-trained models. By generating a series of questions based on the metadata and content of the top-retrieved video and answering the question through LLMs, MERLIN iteratively refines query embedding with answer embedding, thereby enhancing semantic relevance of queries and the information contained within the videos. Experimental results on benchmark datasets, including MSR-VTT, MSVD, and ActivityNet, demonstrate the superiority of the retrieval performance (e.g. Recall@K) by showing significant improvement. Specifically, MERLIN boosts text-video retrieval performance (Recall@1) of Google Multimodal Embedding from $44.00$ to $78.00$ on MSR-VTT, from $52.39$ to $77.61$ on MSVD and from $56.58$ to $68.44$ on ActivityNet.
@@
The key contributions of our paper are as follows: (1) Introduction of MERLIN, a novel LLM-based framework for multimodal embedding refinement that addresses discrepancies between user queries and video content by integrating user perspectives. (2) Implementation of an iterative, cost-effective method for refining query embeddings using LLMs, significantly reducing computational demands while improving retrieval accuracy. (3) Presentation of the first retrieval-rerank pipeline in text-video retrieval, enhancing interactivity and context-awareness within multimodal systems. (4) Experimental results shows that MERLIN substantially improves Recall@1 on MSR-VTT, MSVD and ActivityNet, thereby demonstrating notable enhancements in zero-shot text-video retrieval.
@@
Figure 2: An illustration of MERLIN for text-video retrieval. The yellow arrow represents the LLM Questioner returning a question for next round based on metadata of anchor video (Section 3.3). The green arrow represents the human-simulating LLM agent returning an answer based on the “video in mind” through Question Answering module and Aggregation module (Section 3.4). The pink arrow represents MERLIN returning a retrieved video candidates through Multimodal Encoder and Reranker (Section 3.5). The system initially retrieves video candidates $\hat{v}^{0}$ based on the input query text $q$ using a pre-trained multimodal encoder. Using anchor video, LLM Question Generator produces a question $\hat{q}^{1}$ to elicit additional information from the user (Section 3.3). The LLM Agent answers this question based on the “video in mind”, mimicking the human feedback process $\tilde{a}^{1}$. The query and answer embeddings are then gradually integrated for each round. The updated query embedding is used to rerank the video candidates $\hat{v}^{1}$, and the process repeats for multiple rounds.
@@ 2 Related Works
Prior studies have focused on cross-modal interaction, large-scale foundation model training, and probabilistic modeling. In cross-modal interaction Wang et al. (2023); Huang et al. (2023); Jin et al. (2023) have enhanced reasoning abilities by capturing cross-modal similarities at multiple granularity levels, introduced efficient video prompt mechanisms Lester et al. (2021) with minimal trainable parameters, and improved retrieval with strategies like Disentangled Conceptualization and Set-to-Set Alignment. In foundation model training Chen et al. (2024b, 2023); Zhao et al. (2024); Wang et al. (2024a) significant advances have been made with the development of large-scale video and vision-language models leveraging extensive web data, and fine-tuning techniques for better performance on downstream tasks. In probabilistic modeling Hao et al. (2023); Fang et al. (2023); Hao and Zhang (2023), novel alignment methods and modeling of video and text representations as probabilistic distributions have been proposed to improve text-video retrieval accuracy and address domain adaptation challenges.
Concurrent to prior studies and our work, Levy et al. (2023) proposed a chat-based image retrieval system (ChatIR) that interacts with users through conversation to gather additional information beyond the initial query, aiming to better understand and clarify the user’s search intent. Different from ChatIR, MERLIN incorporates frame-level answer generation tailored to the specific requirements of text-video retrieval, employing a training-free approach. Furthermore, inspired by Composed Image Retrieval Liu et al. (2021); Jang et al. (2024), we iteratively refine the embedding by employing spherical linear interpolation, instead iteratively concatenate question and answer pair and feed into retrieval model. Lastly, we handle both multi-modality data simultaneously, meaning that our generation result would be more likely aligned to user’s search intent. This iterative refinement process mirrors human tendencies to continuously improve their queries based on interactive feedback, akin to strategies seen in feedback-based refinement in textual content. This approach is supported by the growing application of reinforcement learning, which has been increasingly utilized to enhance the quality of generated content through both reference-based and reference-independent feedback mechanisms Böhm et al. (2019); Stiennon et al. (2020); Ziegler et al. (2019); Wu et al. (2020); Ouyang et al. (2022); Glaese et al. (2022); Akyürek et al. (2023); Madaan et al. (2023); Lee et al. (2024); Liang et al. (2024).
@@ 3 Multimodal Embedding Refinement via LLM based Iterative Navigation
3.1 Retrieval Scenario
Figure 3: An article of scenario of Google Lens.
In this work, we consider a scenario where a user aims to find a specific video in a search engine, having a firm intent and precise specifications in mind for the desired video. We refer to this desired video as the “video in mind” – an imagined video that the user seeks to retrieve, encompassing various attributes such as visual content, context, and temporal dynamics. However, expressing the full specifications of this “video in mind” through a single query often leads to a discrepancy between the user’s intent and the retrieved results, hindering the search experience.
We incorporate this “video in mind” scenario into the text-video retrieval task, where the goal is to retrieve the most relevant video from a collection based on the user’s textual query. Inspired by the success of multimodal systems like Google Lens [^2] https://lens.google/, which effectively handle and integrate both image and text inputs, we propose MERLIN, an interactive pipeline that dynamically elicits and incorporates user feedback to bridge the gap between the user’s search intent and the retrieved video candidates. MERLIN leverages the power of LLMs to simulate a human-like feedback process, where a “human-simulating LLM agent” engages in a multi-turn conversation with the user, asking clarifying questions and providing answers based on the “video in mind” – the imagined video that the agent has in mind based on the user’s initial query and subsequent feedback.
@@ 3.2 Background
$\mathbf{e}_{q}$ $=f_{\text{enc}}(q)\in\mathbb{R}^{d_{q}}$
$\mathbf{e}_{v}$ $=f_{\text{enc}}(v)\in\mathbb{R}^{d_{v}},$
where $d_{q}$ and $d_{v}$ denote the dimension of query and video embeddings, respectively. The goal of text-video retrieval is to search the most relevant videos $\hat{v}$’s from a collection of videos $\mathcal{V}$ given a query text $q$ as follows:
@@
where $\textsc{sim}(\cdot)$ is a similarity function (e.g., cosine distance, etc). Additionally, our system utilizes two key components: $\mathcal{M}$ and $\mathcal{T}$. Here, $\mathcal{M}$ represents the LLMs and template function $\mathcal{T}$ applies a pre-defined template to inputs [^3] Note that $\mathcal{M}$ is used interchangeably to indicate both a Large Language Model (LLM) and a Multimodal Large Language Model (MLLM)., [^4] Here, subscripts have been omitted for simplicity. However, subscripts are employed in the equations for each specific module (e.g., $\mathcal{M}_{\mathsf{question}}$). In addition, the pre-defined template is presented in Appendix due to the limited space.. Based on this background, we would like to introduce LLM-based iterative navigation, involving multiple rounds of feedback learning and reranking, leading to better performance and interpretability.
@@ 3.3 Question Generation
Suppose that we have retrieved candidates $\hat{v}^{r}_{k}$ where $r$ and $k$ indicate the round and the index of the retrieved top $K$ candidates, respectively. We choose $\hat{v}^{r}_{0}$ as an anchor candidate and generate the question with $\mathcal{M}_{\mathsf{question}}$ as follows [^5] Note that we use the caption from metadata of $\hat{v}^{r}_{0}$ and assume that each video consists of $N$ frames.:
$\hat{q}^{r}=\mathcal{M}_{\mathsf{question}}\big{(}\mathcal{T}_{\mathsf{question}}(\hat{v}^{r}_{0})\big{)}$ (2)
Intuitively, top-ranked candidate is more likely to align with the user’s query. This implies that assessing retrieved candidates with question generated from $\hat{v}^{r}_{0}$ using LLMs would enhance retrieval performance and interpretability.
@@ 3.4 Human-Simulating Agent
Our underlying assumption is mitigating the discrepancy between user queries and the information within the videos would be helpful for the better retrieval performance. To this end, we answer the question $\hat{q}^{r}$ with video $\hat{v}^{r}$ which consists of $N$ frames sampled per second as follows:
$\big{[}\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]}=\mathcal{M}_{\mathsf{answer}}\big{(}\mathcal{T}_{\mathsf{answer}}(\hat{q}^{r}),\hat{v}^{r}_{0}\big{)}$ (3)
It is worth noting that in a real-world scenario, $\mathcal{M}_{\mathsf{answer}}$ could be replaced by a human to consider the user perspective better. Additionally, using $N$ frames allows us to efficiently handle the temporal information inherent to video, capturing the dynamic aspects of the content. This approach enhances our ability to provide a more comprehensive understanding and alignment with the user’s query.
@@
$\tilde{a}^{r}=\mathcal{M}_{\mathsf{aggr}}\Big{(}\mathcal{T}_{\mathsf{aggr}}\big{(}[\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}]\big{)}\Big{)}$ (4)

### v2

@@ Abstract
The rapid expansion of multimedia content has made it increasingly challenging to retrieve relevant videos from large collections accurately. Recent advancements in text-video retrieval have focused on cross-modal interactions, large-scale foundation model training, and probabilistic modeling, yet often neglect the crucial user perspective, leading to discrepancies between user queries and the content retrieved. To address this, we introduce MERLIN (Multimodal Embedding Refinement via LLM-based Iterative Navigation), a novel training-free pipeline that leverages Large Language Models (LLMs) for iterative feedback learning. MERLIN refines query embeddings from a user perspective, enhancing alignment between queries and video content through a dynamic question answering process. Experimental results on datasets like MSR-VTT, MSVD, and ActivityNet demonstrate that MERLIN substantially improves R@1, outperforming existing systems and confirming the benefits of integrating LLMs into multimodal retrieval systems for more responsive and context-aware multimedia retrieval [^1] https://github.com/dhk1349/MERLIN_text_to_video_search.git.
@@
Donghoon Han1111, Eunhwan Park1111, Gisang Lee2, 3111, Adam Lee4, 5, Nojun Kwak6222 1Buzzni AI Lab, 2KAIST, 3Mathpresso Inc., 4UC Berkeley, 5Fainders AI, 6Seoul National University {owen, jude}@buzzni.com, bobopack@kaist.ac.kr, alee00@berkeley.edu, nojunk@snu.ac.kr
@@ 1 Introduction
Multimedia content has recently grown rapidly in both quantity and quality, making the task of finding relevant videos from vast collections increasingly challenging. While recent studies on text-video retrieval have primarily focused on cross-modal interaction Wang et al. (2023); Huang et al. (2023); Wu et al. (2023); Jin et al. (2023), large-scale foundation model training Chen et al. (2024b, 2023); Zhao et al. (2024); Wang et al. (2024a) and probabilistic modeling Hao et al. (2023); Fang et al. (2023); Hao and Zhang (2023), there remains a notable lack of consideration for the discrepancy in text-video retrieval. For instance, as illustrated in Figure 1, the video caption “a baby playing with a cat’s tail” fails to fully capture the additional context of a playful interaction between the baby and the cat. In real-world scenarios, such discrepancies often arise because users tend to submit succinct queries that do not capture the full context of the videos related to their search intent. Consequently, this mismatch can lead to unsatisfactory retrieval performance. Moreover, neglecting the user perspective makes users refine their natural language query multiple times to fully reflect their search intent. This degrades the quality of user experience and makes it difficult to understand the search intent, leading to a discrepancy between user queries and the information within the retrieved videos.
@@
Figure 1: An illustration of the discrepancy between the video caption which could be treated as a user query and the video from MSR-VTT dataset. Blue indicates the details that can be observed statically within the video frame, while red reflects the information that can be obtained temporally across multiple frames.
@@
To address this issue, we introduce MERLIN (Multimodal Embedding Refinement via LLM-based Iterative Navigation), a novel training-free and iterative feedback learning pipeline that leverages the power of Large Language Models (LLMs) to augment queries based on the user perspective, thereby mitigating the aforementioned discrepancies and significantly improving the text-video retrieval performance. Inspired by human problem-solving and cognitive feedback mechanisms Flower and Hayes (1981); Doherty and Balzer (1988), we employ an interactive and iterative feedback learning Böhm et al. (2019); Stiennon et al. (2020); Ziegler et al. (2019); Wu et al. (2020); Ouyang et al. (2022); Glaese et al. (2022); Akyürek et al. (2023); Madaan et al. (2023); Lee et al. (2024a); Liang et al. (2024) consisting of a question answering process that iteratively refines query embeddings for text-video retrieval. Moreover, to our best knowledge, MERLIN presents the first implementation of a retrieval–rerank pipeline in the domain of text-video retrieval, establishing a novel framework that prioritizes user intention and interaction in refining search results.
The primary strength of MERLIN lies in its capability to iteratively adapt and refine query embeddings without necessitating the costly re-training of pre-trained models. As shown in Figure 2, when a user submits a query, MERLIN generates questions based on the metadata of the retrieved video candidates and presents these questions to the user. By gathering additional information from the user’s responses, MERLIN refines the embeddings to improve retrieval accuracy, thereby helping users find “video in mind”222“video in mind” refers to the specific video users are looking for or have in mind during the search process..
Experimental results on benchmark datasets, including MSR-VTT, MSVD, and ActivityNet, demonstrate the superiority of the retrieval performance (e.g. R@K) by showing significant improvement. Specifically, MERLIN boosts text-video retrieval performance (R@1) of Google Multimodal Embedding from $44.00$ to $78.00$ on MSR-VTT, from $52.39$ to $77.61$ on MSVD and from $56.58$ to $68.44$ on ActivityNet.
@@
The key contributions of our paper are as follows: (1) Introduction of MERLIN, a novel LLM-based framework for multimodal embedding refinement that addresses discrepancies between user queries and video content by integrating user perspectives. (2) Implementation of an iterative, cost-effective method for refining query embeddings using LLMs, significantly reducing computational demands while improving retrieval accuracy. (3) Presentation of the first retrieval-rerank pipeline in text-video retrieval, enhancing interactivity and context-awareness within multimodal systems. (4) Experimental results shows that MERLIN substantially improves R@1 on MSR-VTT, MSVD and ActivityNet, thereby demonstrating notable enhancements in zero-shot text-video retrieval.
@@
Figure 2: An illustration of MERLIN for text-video retrieval. The yellow arrow represents the LLM Questioner returning a question for next round based on metadata of anchor video (Section 3.2). The green arrow represents the human-simulating LLM agent returning an answer based on the “video in mind” through Question Answering module and Aggregation module (Section 3.3). The pink arrow represents MERLIN returning a retrieved video candidates through Multimodal Encoder and Reranker (Section 3.4). The system initially retrieves video candidates $\hat{v}^{0}$ based on the input query text $q$ using a pre-trained multimodal encoder. Using this anchor video, LLM Question Generator produces a question $\hat{q}^{1}$ to elicit additional information from the user (Section 3.2). The LLM Agent answers this question based on the “video in mind”, mimicking the human feedback process $\tilde{a}^{1}$. The query and answer embeddings are then gradually integrated for each round. The updated query embedding is used to rerank the video candidates $\hat{v}^{1}$, and the process repeats for multiple rounds.
@@ 2 Related Works
Prior studies have focused on cross-modal interaction, large-scale foundation model training, and probabilistic modeling. In cross-modal interaction Wang et al. (2023); Huang et al. (2023); Jin et al. (2023) have enhanced reasoning abilities by capturing cross-modal similarities at multiple granularity levels, introduced efficient video prompt mechanisms Lester et al. (2021) with minimal trainable parameters, and improved retrieval with strategies like Disentangled Conceptualization and Set-to-Set Alignment. In foundation model training Chen et al. (2024b, 2023); Zhao et al. (2024); Wang et al. (2024a), significant advances have been made with the development of large-scale video and vision-language models leveraging extensive web data, and fine-tuning techniques for better performance on downstream tasks. In probabilistic modeling Hao et al. (2023); Fang et al. (2023); Hao and Zhang (2023), novel alignment methods and modeling of video and text representations as probabilistic distributions have been proposed to improve text-video retrieval accuracy and addressed domain adaptation challenges.
Concurrent to prior studies, Levy et al. (2023) proposed a chat-based image retrieval system (ChatIR) that interacts with users through conversation to gather additional information beyond the initial query, aiming to better understand and clarify the user’s search intent. Following from ChatIR, Lee et al. (2024b) proposed the plug-and-play interactive text-to-image retrieval system. Different from ChatIR, our MERLIN incorporates frame-level answer generation tailored to the specific requirements of text-video retrieval, employing a training-free approach. Furthermore, inspired by Composed Image Retrieval Liu et al. (2021); Jang et al. (2024), we iteratively refine the embedding by employing spherical linear interpolation, instead of iteratively concatenating question and answer pair and feeding into the retrieval model. Lastly, we handle both multi-modality data simultaneously, meaning that our generation result would be more likely aligned to the user’s search intent. This iterative refinement process mirrors human tendencies to continuously improve their queries based on interactive feedback, akin to strategies seen in feedback-based refinement in textual content. This approach is supported by the growing application of reinforcement learning, which has been increasingly utilized to enhance the quality of generated content through both reference-based and reference-independent feedback mechanisms Böhm et al. (2019); Stiennon et al. (2020); Ziegler et al. (2019); Wu et al. (2020); Ouyang et al. (2022); Glaese et al. (2022); Akyürek et al. (2023); Madaan et al. (2023); Lee et al. (2024a); Liang et al. (2024).
@@ 3.1 Background
1:encoder $f_{\text{enc}}()$, user query $q\in\mathcal{Q}$, video $v\in\mathcal{V}$, total question answer round $R$, retrieved top-$k$ videos at round $r$ $\hat{v}^{r}$, $i$-th candidate among top-$k$ videos at round $r$ $\hat{v}^{r}_{i}$, $v^{m}$ a video that user is looking for
2:Encode $\mathbf{e}_{q}=f_{\text{enc}}(q)$ given user query $q$
3:Encode $\mathbf{e}_{v}=f_{\text{enc}}(v)$ given video $v$
5:Initialize message list $m=[]$
6:for $r=1$ to $R$ do
7:     Append metadata of $\hat{v}^{r-1}_{0}$ to $m$
9:     Append $\hat{q}^{r}$ to $m$
10:     Generate frame-level answers $\big{[}\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]}=\mathcal{M}_{\mathsf{answer}}(\hat{q}^{r}:v^{m})$ (Equation 3)
12:     Encode $\mathbf{e}_{A^{r}}=f_{\text{enc}}(A^{r})$
13:     Refine embedding $e=\textsc{Refine}(e_{q},\cdots,e_{A^{r}})$ (Equation 6)
15:end for
16:return Reranked retrieved videos $\hat{v}^{r}_{k}$
@@
$\mathbf{e}_{q}$ $=f_{\text{enc}}(q)\in\mathbb{R}^{d}$
$\mathbf{e}_{v}$ $=f_{\text{enc}}(v)\in\mathbb{R}^{d},$
where $d$ denotes the dimension of embedding. The goal of text-video retrieval is to search the most relevant videos $\hat{v}$’s from a collection of videos $\mathcal{V}$ given a query text $q$ as follows:
@@
where $\textsc{sim}(\cdot)$ is a similarity function (e.g., cosine distance, etc). Additionally, our system utilizes two key components: $\mathcal{M}$ and $\mathcal{T}$. Here, $\mathcal{M}$ represents the LLMs and template function $\mathcal{T}$ applies a pre-defined template to inputs [^3] Note that $\mathcal{M}$ is used interchangeably to indicate both a Large Language Model (LLM) and a Large Multimodal Model (LMM)., [^4] Here, subscripts have been omitted for simplicity. However, subscripts are employed in the equations for each specific module (e.g., $\mathcal{M}_{\mathsf{question}}$). In addition, the pre-defined template is presented in Appendix due to the limited space.. Based on this background, we would like to introduce LLM-based iterative navigation, involving multiple rounds of feedback learning and reranking, leading to better performance and interpretability.
@@ 3.2 Question Generation
Suppose that we have retrieved candidates $\hat{v}^{r}_{k}$ where $r$ and $k$ indicate the round and the index of the retrieved top $K$ candidates, respectively. We choose $\hat{v}^{r-1}_{0}$ as an anchor candidate and generate the question with $\mathcal{M}_{\mathsf{question}}$ as follows [^5] Note that we use the caption from metadata of $\hat{v}^{r}_{0}$ and assume that each video consists of $N$ frames.:
$\hat{q}^{r}=\mathcal{M}_{\mathsf{question}}\big{(}\mathcal{T}_{\mathsf{question}}(\hat{v}^{r-1}_{0})\big{)}.$ (2)
Intuitively, top-ranked candidate is more likely to align with the user’s query. This implies that assessing retrieved candidates with question generated from $\hat{v}^{r-1}_{0}$ using LLMs would enhance retrieval performance and interpretability.
@@ 3.3 Human-Simulating Agent
Our underlying assumption is mitigating the discrepancy between user queries and the information within the videos would be helpful for the better retrieval performance.
To this end, a human-simulating agent answers the question $\hat{q}^{r}$ with video in mind $v^{m}$, which consists of N frames sampled per second as follows. In this process, we assume a user searching for a specific video, and create a human-simulating agent to mimic the behavior of that user. The agent generates responses by referencing both the video in mind $v^{m}$ (the video the user is looking for) and the questions generated by MERLIN as following:
$\big{[}\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]}=\mathcal{M}_{\mathsf{answer}}\big{(}\mathcal{T}_{\mathsf{answer}}(\hat{q}^{r}),v^{m}\big{)}$ (3)
It is worth noting that in a real-world scenario, $\mathcal{M}_{\mathsf{answer}}$ could be replaced by a human. Additionally, using $N$ frames allows us to efficiently handle the temporal information inherent to video, capturing the dynamic aspects of the content. This approach enhances our ability to provide a more comprehensive understanding and alignment with the user’s query.
@@
$\tilde{a}^{r}=\mathcal{M}_{\mathsf{aggr}}\Big{(}\mathcal{T}_{\mathsf{aggr}}\big{(}[\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}]\big{)}\Big{)}.$ (4)